OpenStudy (anonymous):

What is the derivative of: sqrt((3x-1)/(2x+3))

OpenStudy (anonymous):

\[\sqrt{(3x-1) / (2x+3)}\] I got: \[15/(2(2x+3)^{2}) * \sqrt{(2x+3)/(3x-1)}\]

OpenStudy (anonymous):

oh thats different than what i thought it was... give me a minute lol

OpenStudy (anonymous):

\[f \prime(x)= (1/2)((3x-1)/(2x+3))^{-1/2}*((3)(2x+3) - (3x-1)(2)/(2x+3)^{2})\]\[((3x-1)/(2x+3)^{2})\]

OpenStudy (anonymous):

i would use the quotient rule with this one and break it down into the equation of (3x-1)^(1/2)/(2x+3)^(1/2) then use the quotient rule and power rule to find the answer. my laptop is about to die so i can write my answer in about 30 min.

OpenStudy (anonymous):

\[=(1/2)((2x+3)/(3x-1))^{1/2} *( (6x+9-6x+6)/(2x+3)^{2})\] \[=(1/2)((2x+3)/(3x-1))^{1/2} *( (15)/(2x+3)^{2})\] \[=((15)/(2(2x+3)^{2}))*\sqrt{(2x+3)/(3x-1)}\]